The perimeter of a right, isosceles triangle is 2 + square root (2)
Find its area. Express your answer as a common fraction.

Answer by edjones(8007)

a^2+b^=c^2
if a=1 and b=1 then c=sqrt(2)
So the perimeter is 2+sqrt(2)
So A=.5ab=.5*1*1=.5=1/2 units^2
.
Ed